### v1

@@ 1 Introduction
The anomalies of time-reversal symmetry in gauge theories with charged fermions have been studied extensively in the literature (e.g. [6, 7, 2, 8]). However, there were several open problems:
@@
In this note we will demonstrate a method of computing the anomalies that address these issues, using gauge theories with $SU(N),Sp(N),U(1),SO(N)$ gauge groups and fermions in the fundamental and rank-two representations. The symmetries discussed here are invertible symmetries. In a upcoming paper we will discuss non-invertible symmetries. There are various proposed dynamics of these gauge theories, many of them involves Goldstone bosons, e.g. [15, 16]. It will be interesting to constrain the dynamics using anomalies in future work.
@@ 1.3 Summary of results
  • By stacking $SU(16n+8)$ gauge theory with rank-2 representation and e.g. $SU(2)$ gauge theory with two fundamental fermions, we obtain bosonic theories with $\alpha=1,\beta=0$. In other words, the combined bosonic theory has a multiple of 16 Majorana fermions and $\alpha=0$, but the time-reversal anomaly $w_{2}^{2}$ is nontrivial $\beta=1$.
@@ 5.3 Anomaly of different anti-unitary symmetries $T^{\prime}=TU$
We can combine the time-reversal symmetry $T$ that commutes with the gauge group with another unitary symmetry $U$ to define a new time-reversal symmetry $T^{\prime}=TU$. The anomaly of $T^{\prime}$ can be derived from that of $T,U$. Let us illustrate this procedure using the example of QED3 with $N_{f}=4$ Dirac fermions. The theory has $T^{2}=1$ time-reversal symmetry, and is proposed to describe Dirac spin liquid in certain Heisenberg antiferromagnet (e.g. [34]).444 We thank Meng Cheng for bring up this example. The theory is also used to describe an “unnecessary” quantum critical point on the lattice [35].
@@ 5.3.2 New $T^{\prime}$ that combines flavor charge conjugation symmetry
In [34] the magnetic charge is odd under time-reversal, while here the magnetic charge is even under time-reversal. In addition, the time-reversal symmetry flips the sign of 3 out of the 6 monopoles that form 2-index antisymmetric representation of $SU(4)$ flavor symmetry. The 6 monopoles carry $SU(4)$ fundamental indices ${\cal M}_{1,2},{\cal M}_{1,3},{\cal M}_{1,3},{\cal M}_{2,3},{\cal M}_{2,4},{\cal M}_{3,4}$. Thus the internal symmetry is the flavor charge conjugation symmetry ${\cal C}_{f}$ that flips the sign of index $1$. In other words, the time-reversal symmetry is $T^{\prime}=T{\cal C}{\cal C}_{f}$. By combining the previous two cases, one finds both $w_{2}^{2}$ and $w_{1}^{4}$ anomalies are absent, in agreement with [34].
@@ 6.3 Discrete gauge theory with fermions
Here we provide a universal explanation for the absence of $w_{2}^{2}$ anomaly in gauge theory without one-form symmetry in section 2-4.
To compute the anomalies, we can introduce adjoint scalars that are even under time-reversal symmetry.888 This is the case for the time-reversal symmetry $T$ that commutes with the gauge group. The theory is still bosonic, since the fermion parity, which is in the center of the gauge group, does not act on the adjoint scalars. We do not turn on any Yukawa couplings. When the scalars do not condense, the theory with scalars flow to the original gauge theories with charged fermions. When the scalars condense, we can use sufficiently number of scalars and generic potential such that the gauge group is broken to the center. Then the anomaly can be computed in Abelian gauge theories coupled to fermions. The Abelian gauge groups can be product of $U(1)$ or $\mathbb{Z}_{2}^{f}$ fermion parity. The former is already discussed in section 5 where it is shown that the $w_{2}^{2}$ anomaly is absent. Here we will focus on the latter case.
In the microscopic theory without one-form symmetry, we are left with $\mathbb{Z}_{2}^{f}$ gauge theory coupled to $N_{f}$ Dirac fermions with even $N_{f}$, where $f$ superscript indicates that it is identified with $(-1)^{F}$ fermion parity. To preserve time-reversal symmetry, we need to add the counterterm
$(\mathbb{Z}_{2})^{f}_{N_{f}}+N_{f}\text{CS}_{\text{grav}}~{}.$ (6.17)
@@ Acknowledgement
We thank Ryan Thorngren, Cenke Xu and Yi-Zhuang You for related discussions. We thank Meng Cheng for discussions and comments on a draft. We thank Thomas Dumitrescu, Jaume Gomis and Ho tat Lam for comments on a draft. The work was supported by Simons Collaboration of Global Categorical Symmetry, Department of Mathematics King’s College London, and also supported in part by grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP). We thank Kavli Institute for Theoretical Physics for hosting the program “Correlated Gapless Quantum Matter” in 2024, during which part of the work is completed.
@@ Appendix A Useful Mathematical Identities for Characteristic Classes
  • Wu formula. On a general closed $D$-manifold and $\mathbb{Z}_{2}$ $n$-cocycle $x_{n}$, $Sq^{D-n}x_{n}=x_{n}\cup_{2n-D}x_{n}=v_{D-n}\cup x_{n}$, where $v_{D-n}$ is the $(D-n)$th Wu class (see e.g. [29]). For $D=4$ and $n=2$, this is $x_{2}\cup x_{2}=Sq^{2}(x_{2})=(w_{2}+w_{1}^{2})\cup x_{2}$, where we used the second Wu class $v_{2}=w_{2}+w_{1}^{2}$.
@@
    Using $w_{4}+w_{2}^{2}+w_{1}^{4}=0$ mod 2 this becomes

### v2

@@ 1 Introduction
The anomalies of time-reversal symmetry in gauge theories with charged fermions have been studied extensively in the literature (e.g. [6, 7, 2, 8, 9, 10, 11]). However, there were several open problems:
@@
In this note we will demonstrate a method of computing the anomalies that address these issues, using gauge theories with $SU(N),Sp(N),U(1),SO(N)$ gauge groups and fermions in the fundamental and rank-two representations. The symmetries discussed here are invertible symmetries. In a upcoming paper we will discuss non-invertible symmetries. There are various proposed dynamics of these gauge theories, many of them involve Goldstone bosons, e.g. [18, 19]. It will be interesting to constrain the dynamics using anomalies in future work.
@@ 1.3 Summary of results
  • By stacking $SU(16n+8)$ gauge theory with rank-2 representation and e.g. $SU(2)$ gauge theory with two fundamental fermions, we obtain bosonic theories with $\alpha=1,\beta=0$. The combined bosonic theory has a multiple of 16 Majorana fermions and $\beta=0$, but the time-reversal anomaly $w_{2}^{2}$ is nontrivial $\alpha=1$. The theories have $c_{-}=4$ mod 8 anomaly, which can be different from free massless fermions contribution.
@@ 5.3 Anomaly of different anti-unitary symmetries $T^{\prime}=TU$
We can combine the time-reversal symmetry $T$ that commutes with the gauge group with another unitary symmetry $U$ to define a new time-reversal symmetry $T^{\prime}=TU$. The anomaly of $T^{\prime}$ can be derived from that of $T,U$. Let us illustrate this procedure using the example of QED3 with $N_{f}=4$ Dirac fermions. The theory has $T^{2}=1$ time-reversal symmetry, and is proposed to describe Dirac spin liquid in certain Heisenberg antiferromagnet (e.g. [8]).444 We thank Meng Cheng for bringing up this example. The anomalies of $N_{f}=4$ has been computed in [9, 10], where the computation considers the background gauge field of internal symmetry to be $SU(4)\times U(1)$ without $\mathbb{Z}_{2}$ quotient, and the $\mathbb{Z}_{2}$ charge conjugation. We remark that the theory is also used to describe an “unnecessary” quantum critical point on the lattice [37].
@@ 5.3.2 New $T^{\prime}$ that combines flavor charge conjugation symmetry
In [8] the magnetic charge is odd under time-reversal, while here the magnetic charge is even under time-reversal. In addition, the time-reversal symmetry flips the sign of 3 out of the 6 monopoles that form 2-index antisymmetric representation of $SU(4)$ flavor symmetry. The 6 monopoles carry $SU(4)$ fundamental indices ${\cal M}_{1,2},{\cal M}_{1,3},{\cal M}_{1,3},{\cal M}_{2,3},{\cal M}_{2,4},{\cal M}_{3,4}$. Thus the internal symmetry is the flavor charge conjugation symmetry ${\cal C}_{f}$ that flips the sign of index $1$. In other words, the time-reversal symmetry is $T^{\prime}=T{\cal C}{\cal C}_{f}$. By combining the previous two cases, one finds both $w_{2}^{2}$ and $w_{1}^{4}$ anomalies are absent, in agreement with [8, 9, 10].
@@ 6.3 Discrete gauge theory with fermions
Here we provide a universal explanation for the absence of $w_{2}^{2}$ anomaly in gauge theory without one-form symmetry in section 2-4. For simplicity, we will assume the total number of Dirac fermions is even, which is the case in all our examples in section 2-4.
To compute the anomalies, we can introduce adjoint scalars $\phi^{A}$ that are even under time-reversal symmetry.888 This is the case for the time-reversal symmetry $T$ that commutes with the gauge group. For instance, the spatial components of the gauge field, which is also in the adjoint representation, transform as $a_{i}(t,\vec{x})\rightarrow a_{i}(-t,\vec{x})$. The theory is still bosonic, since the fermion parity, which is in the center of the gauge group, does not act on the adjoint scalars. We do not turn on any Yukawa couplings. When the scalars do not condense, the theory with scalars flow to the original gauge theories with charged fermions. When the scalars condense, we can use sufficiently number of scalars and generic potential such that the gauge group is broken to the center. Then the anomaly can be computed in Abelian gauge theories coupled to fermions. The Abelian gauge groups can be product of $U(1)$ or $\mathbb{Z}_{2k}\supset\mathbb{Z}_{2}^{f}$ fermion parity. The former is already discussed in section 5 where it is shown that the $w_{2}^{2}$ anomaly is absent. Here we will focus on the latter case. Next, We further add charge-two Higgs scalars $\Phi$ with real condensate and no Yukawa couplings such that the gauge group is further broken down to $\mathbb{Z}_{2}^{f}$ fermion parity, while preserving the original time-reversal symmetry $\Phi(t,\vec{x})\rightarrow\Phi^{*}(-t,\vec{x})$.
The microscopic theory without one-form symmetry flows to $\mathbb{Z}_{2}^{f}$ gauge theory coupled to $N_{f}$ Dirac fermions with even $N_{f}$. To preserve time-reversal symmetry, we need to add the counterterm
$(\mathbb{Z}_{2}^{f})_{N_{f}}+N_{f}\text{CS}_{\text{grav}}~{}.$ (6.17)
@@ Acknowledgement
We thank Ryan Thorngren, Cenke Xu and Yi-Zhuang You for related discussions. We thank Meng Cheng and Chong Wang for discussions and comments on a draft. We thank Thomas Dumitrescu, Jaume Gomis and Ho tat Lam for comments on a draft. The work was supported by Simons Collaboration of Global Categorical Symmetry, Department of Mathematics King’s College London, and also supported in part by grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP). We thank Kavli Institute for Theoretical Physics for hosting the program “Correlated Gapless Quantum Matter” in 2024, during which part of the work is completed.
@@ Appendix A Useful Mathematical Identities for Characteristic Classes
  • Wu formula. On a general closed $D$-manifold and $\mathbb{Z}_{2}$ $n$-cocycle $x_{n}$, $Sq^{D-n}x_{n}=x_{n}\cup_{2n-D}x_{n}=v_{D-n}\cup x_{n}$ mod 2, where $v_{D-n}$ is the $(D-n)$th Wu class (see e.g. [32]). For $D=4$ and $n=2$, this is $x_{2}\cup x_{2}=Sq^{2}(x_{2})=(w_{2}+w_{1}^{2})\cup x_{2}$ mod 2, where we used the second Wu class $v_{2}=w_{2}+w_{1}^{2}$.
@@
    Using $w_{4}+w_{2}^{2}+w_{1}^{4}=0$ mod 2 on four-manifolds, this becomes